Purple Company has $200,000 in net income for 2018 before deducting any compensation or other payment to its sole owner, Kirsten. Kirsten is single and she claims the $12,000 standard deduction for 2018. Purple Company is Kirsten's only source of income. Ignoring any employment tax considerations, compute Kirsten's after-tax income for each of the following situations. Assume the corporate tax rate is 21%. When required, carryout intermediate tax computations to the nearest cent and then round your final tax liability to the nearest dollar. a. If Purple Company is a proprietorship and Kirsten withdraws $50,000 from the business during the year; Kirsten claims a $40,000 deduction for qualified business income ($200,000 × 20%). Kirsten's taxable income is $_____, and her after-tax income is $_________. b. Purple Company is a C corporation and the corporation pays out all of its after-tax income as a dividend to Kirsten. Note: Individual taxpayers received preferential treatment regarding the taxation of qualified dividends (0%,15%,20%). For single taxpayers, the 0 percent rate applies to the first $38,600 of taxable income. Purple Corporation's after-tax income is $______ and Kristen's after tax income is $_______. c. Purple Company is a C corporation and the corporation pays Kirsten a salary of $158,000. Kirsten's after-tax income is $_________.
Answer #1

Ans) IN CASE ONE

Kristen's Taxable Income-

Total Income from Business- $200,000

less; Standard Deduction- $12000

Taxable Income- $188,000

Kristen's After Tax Income-  

Taxable Income $188000

less; tax on $188,000 $42738.25

  (18713.75+ 28% of amount over

$91900)

After Tax Income- $157261.75

IN CASE TWO

Tax Deducted on total earnings- $22250+ 39% of earnings over $100000= $22250+$39000= $61250

After tax income for the business- $200000-$61250= $138750

Therefore Kristen's taxable income- $138750-$12000=$126750

Payable tax= ($38600*0%)+ 15%($126750-$38600)=$13222.50

Kristen's Income after tax ($138750-$13222.50) = $125527.50

IN CASE THREE

Tax Liability= $18713.75+ 28% of amount above $91900(i.e $126750-$91900)

= $18713.75+ 28% $34850=$28471.75

After Tax Income= $126750-$28471.75= $98278.75